In right triangleABC , the length of the hypotenuse is 25 and the lengths of the legs are x and 3x + 3. What is
the perimeter of the triangle?

Found 3 solutions by solver91311, stanbon, mananth:
Answer by solver91311(24713)



Since it is a right triangle, call upon our old friend Pythagoras to help:



Solve for , then calculate , then add the measures of the three sides to get the perimeter.

John

My calculator said it, I believe it, that settles it

Answer by stanbon(75887)

In right triangle ABC , the length of the hypotenuse is 25 and the lengths of the legs are x and 3x + 3. What is
the perimeter of the triangle?
-----
Solve for "x"::
x^2 + (3x+3)^2 = 25^2
------
x^2 + 9x^2 + 18x + 9 - 625 = 0
-----
10x^2 + 18x - 616 = 0
---
5x^2 + 9x - 308 = 0
---
(x-7)(5x+44) = 0
---
Positive solution::
x = 7
3x+3 = 24
----
Ans: Perimeter = 25 + 24 + 7 = 56
----------
Cheers,
Stan H.
--------------

Answer by mananth(16946)


Using Pythagoras theorem
x^2+(3x+3)^2 = 625

x^2+9x^2+18x +9 =625
10x^2+18x -616=0
/2
5x^2+9x-308=0
5x^2-35x+44x-308=0
5x(x-7)+44(x-7)=0
(5x+4)(x-7)=0
5x+4 = 0 OR x-7=0
x= -4/5 Or 7
x cannot be negative
Therefore x =7
Check
x^2= 49
(3x+3)^2 = 576
576 + 49 = 625
x= 7
(3x+3)= 24
Hypotenuse = 25
Perimeter = 7+24+25 = 56